OpenStudy (anonymous):

Simplify the product. 5b(6b - 10)

OpenStudy (mathstudent55):

Use the distributive property to distribute the 5b by the 6b and -10 terms. Here are two examples of the distributive property: a(b - c) = ab - ac 6(3b - 8) = 18b - 48

OpenStudy (anonymous):

5b2 + 6b - 10

OpenStudy (mathstudent55):

You need to multiply 5b by 6b, and multiply 5b by - 10

OpenStudy (mathstudent55):

5b(6b - 10) = 5b*6b - 5b*10 What is 5b*6b = What is 5b*10 =

OpenStudy (mathstudent55):

The answer will have only two terms, not three.

OpenStudy (anonymous):

11b - 10

OpenStudy (mathstudent55):

5b * 6b = 5 * 6 * b * b = 30b^2 5b * 10 = 5 * 10 * b = 50b 5b(6b - 10) = 5b*6b - 5b*10 = 30b^2 - 50b

OpenStudy (mathstudent55):

11b = 5b + 6b. You added 5b and 6b. You need to multiply them together: 5b * 6b = 30b^2 Also, the -10 inside the parenteses must also be multiplied by 5b, so 5b * 10 = 50b 5b(6b - 10) = 30b^2 - 50b

OpenStudy (anonymous):

30b2 - 50b

OpenStudy (mathstudent55):

correct